OpenStudy (anonymous):

Solve x^2+10x+14=0 by completing the square?

OpenStudy (anonymous):

It cannot be done by factoring

OpenStudy (anonymous):

subtract \(14\) and get \[x^2+10x=-14\]

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

and then divide the coefficient 10 by 2?

OpenStudy (anonymous):

half of 10 is 5, 5 squared is 25, next line should be \[(x+5)^2=-14+25\] and to from there

OpenStudy (anonymous):

yes, what you said

OpenStudy (anonymous):

x+5 = sqrt(11)

OpenStudy (anonymous):

don't forget the \(\pm\)

OpenStudy (anonymous):

right what is the alt code for that?

OpenStudy (anonymous):

\pm

OpenStudy (anonymous):

\pm

OpenStudy (anonymous):

/pm

OpenStudy (anonymous):

\( first

OpenStudy (anonymous):

and end with \)

OpenStudy (anonymous):

anyways is x = -sqrt(11)-5 x = sqrt(11)-5

OpenStudy (anonymous):

if you want to see any latex code, right click "show math as" then "latex"

OpenStudy (anonymous):

oh /pm\

OpenStudy (anonymous):

\pm/

OpenStudy (anonymous):

\(\pm\)

OpenStudy (anonymous):

right click and you will see it

OpenStudy (anonymous):

right click on what?

OpenStudy (anonymous):

the code

OpenStudy (anonymous):

\(\LaTeX\) right click on the \(\huge \LaTeX\)

OpenStudy (anonymous):

show math as tex commands?

OpenStudy (anonymous):

ok \[x+5=±\sqrt{11}\]

OpenStudy (anonymous):

\[x=-\sqrt{11}-5 , x=\sqrt{11}-5\]

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

traditionally written as \[-5\pm\sqrt{11}\] but it is the same

OpenStudy (anonymous):

OK, thanks :)